OpenStudy (anonymous):

determine the minimum coefficient of friction between the rites and a flat horizontal road that will enable cars to complete a turn of radius 125 m at 80.3 km/h. please show work, im trying to study for a test!

OpenStudy (espex):

Do you have a mass for the car or can you assume one? \[v=\frac{80km}{1h} \times \frac{1h}{3600 \sec} \times \frac{1000m}{1km}=\frac{22.22}{\sec}\] If you assume the mass of a car is 1000kg \[F_r=m\frac{v^2}{r} \rightarrow F_r=1000\frac{(22.22\frac{m}{s})^2}{125m}=3949.83N\] \[F_f=\mu F_n \rightarrow \mu mg \rightarrow \mu=\frac{F_f}{mg} \rightarrow \frac{3949.83N}{1000kg \times 9.8\frac{m}{s^2}}=0.403\]